OpenStudy (anonymous):

5 litres of glycerin contains 20% adulteration. How much glycerin should be added in order to reduce the adulteration upto 5%?

OpenStudy (anonymous):

@amistre64 pls help..

OpenStudy (amistre64):

does pure glycerin contain adulteration to begin with?

OpenStudy (amistre64):

5(.20) + g(0) = (5+g)(.15) is what im thinking

OpenStudy (anonymous):

@amistre64 How you got (.15)

OpenStudy (anonymous):

@amistre64 How you got (.15)

OpenStudy (anonymous):

@amistre64 How you got (.15) pls come here

OpenStudy (amistre64):

if we reduce the adulteration up to 5%; then 20-5 = 15

OpenStudy (amistre64):

we want a 15% concentration of whatever that is

OpenStudy (amistre64):

5(20%) + g(0%) has to be equal to (5+g) at 15%

OpenStudy (anonymous):

No, we have to bring the adulteration level upto 5%

OpenStudy (amistre64):

you might have to define the terminology for me then :)

OpenStudy (anonymous):

Actually i'm supposed to find the adulteration level upto 5%

OpenStudy (amistre64):

5 litres of glycerin contains 20% adulteration How much glycerin should be added in order to reduce (diminish, lower, decrease) the adulteration up to 5%

OpenStudy (anonymous):

Yes,

OpenStudy (amistre64):

so are we reducing the 20% by 5%, to get to 15%? or are we reducing the 20% all the way down to 5%?

OpenStudy (amistre64):

in either case, the calculation are the same, just the "final" results are different

OpenStudy (anonymous):

are we reducing the 20% all the way down to 5%?

OpenStudy (amistre64):

then: 5 at 20% + g at 0% = (5+g) at 5% 5(.2) = 5(.05) + g(.05) 5(.20-.05) = g(.05) 5(.20-.05)/(.05) = g

OpenStudy (amistre64):

i read it as reducing the 20% down to 15% but the process is the same

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

May be that i get back to you again.

OpenStudy (anonymous):

@amistre64 why 'g at' 0%?

OpenStudy (amistre64):

the glycerin added doesnt come with any percentage of adulteration to assess; i assume its at 0%, since 1% only increase the lot

OpenStudy (amistre64):

spose you have a 20% acid solution, how much 0% water would you add to it to reduce it to 5% (or 15%) ?

OpenStudy (amistre64):

try adding a 100% acid solution to it and see if it reduces :)